show your work
please help me with this problem
5 and 2/3x - 16 = 50

Answer by fractalier(2101)

Okay from
(5 2/3)x - 16 = 50
add 16 first to both sides and get
(5 2/3)x = 66
now convert the fraction into improper form
(17/3)x = 66
now multiply both sides by 3/17 and get
x = 66(3/17)
x = 198/17
or if you need it in mixed number form
x = 11 11/17